# Leveraging Modality Tags for Enhanced Cross-Modal Video Retrieval

### v1

@@ III Modality Auxiliary Concepts for Video Retrieval
III-A Cross-Modal Text-to-Video Retrieval
Given a pair $(v_{i},t_{i})$, where $v_{i}$ represents a video and $t_{i}$ denotes its corresponding caption, the objective of Cross-Modal Text-to-Video Retrieval is to retrieve $v_{i}$ from a large gallery of videos, given the caption $t_{i}$ as query. Typically, models use two projection functions: $f_{v}:v_{i}\longrightarrow\Omega\in\mathbb{R}^{d}$ and $f_{t}:t_{i}\longrightarrow\Omega\in\mathbb{R}^{d}$ to map the respective modalities into a shared d-dimensional latent embedding space, $\Omega$. The aim is to align the representations in this space so that the representation of a video is close to that of its corresponding caption. Following training, the standard inference strategy (IS) embeds a test video gallery and ranks these based on their distance from each query caption. Recent approaches utilise additional inference strategies to improve performance: Querybank Normalisation (QB) [17] and Dual Softmax (DSL) [18] are the most popular. We introduce these here and later show their impact across current SOTA.
The QB strategy [17] was introduced to mitigate the hubness problem of high-dimensional embedding spaces [68], where a small subset of samples tends to appear far more frequently among the k-nearest neighbours queries. To mitigate this, the similarities between embeddings are altered to minimise the influence of hubs. A set of samples (the querybank) is sampled from the queries within the training set and used as a probe to measure the hubness of the gallery at test time. More formally, given the vector of unnormalised similarities at test time, $S(v_{j},t_{i})$ and a probe matrix $P$, in which each row is a probe vector of similarities between the querybank and each element in the gallery, we re-weight the query-gallery similarities: $\eta_{q}=\text{QB}(S(v_{j},t_{i}),P)$, where the QB is the querybank normalisation function using Dynamic Inverted Softmax (DIS), introduced in [17].
The second commonly used inference strategy, the DSL strategy [18], introduces an intrinsic prior of each pair in a batch to correct the similarity matrix and achieves the dual optimal match. In practice, we modify the original $S(v_{j},t_{i})$ by multiplying it with a prior $r_{i,j}$. Therefore, we can define the new similarity matrix as $\hat{S}(v_{j},t_{i})=r_{i,j}S(v_{j},t_{i})$, where the prior is defined as $r_{i,j}=\frac{exp(\tau_{r}S(v_{i},t_{i}))}{\sum_{j}exp(\tau_{r}S(v_{i},t_{j}))}$, where $\tau_{r}$ is a temperature hyper-parameter to smooth the gradients. While this strategy can be used both in training and inference, it is now regularly used only during inference.
III-B Tag Extraction
We propose to extract tags automatically from either the video $v_{i}$, using a VLM, or the text in the caption $t_{i}$, using an LLM. These tags are word-level representations of common objects, actions, or general ideas present in the video or the caption. They can add additional useful information to retrieve the correct video given a text query as shown in Fig. 2. For instance, given the caption: a commercial for the Mazda 3 the car sliding around a corner, the general tags estimated from this caption are: product showcase, advertising technique, automotive marketing, which reflect the commercial. These words are abstract terms that go beyond the exact caption but can help the model to understand the specific characteristics of this caption better.
@@ III-B Tag Extraction
To extract visual and textual tags, we use a custom prompt (see Sec. A in appx.) to query the most relevant general tags for the input video $v_{i}$ and caption $t_{i}$. We extract tags individually from both modalities so they can be used for training and inference. As the tags are extracted from a single modality, which we call modality-specific tags. We detail how these tags are used within MAC-VR next.
III-C Architecture
@@
We start from a standard Text-Conditioned Video Encoder (T-CVE) before incorporating our proposed tags into each modalities’ latent concepts. These concepts are aligned and pooled to find the similarity between a video and a caption. Our proposed architecture is summarised in Fig. 3.
Text-Conditioned Video Encoder (T-CVE)
@@
Given a caption $t_{i}$, we extract its text representation ${T_{i}\in\mathbb{R}^{d}}$. For the video representation, we first sample $N_{v}$ frames from a video $v_{i}$ and then encode them and aggregate the embedding of all frames to obtain the frame representation $F_{j}$ with $j\in\{1,...,N_{v}\}$. Since captions often describe specific moments, as shown in previous works [69, 3, 26], matching only the relevant frames improves semantic precision and reduces noise. To achieve this, we aggregate the frame representations conditioned on the text. Firstly, we calculate the inner product between the text and the frame representation $F_{j}$ with $j\in\{1,...,N_{v}\}$:
@@
To utilise the visual and textual tags extracted from foundational models in Sec III-B, we first randomly pick $N$ visual and textual tags (per modality) during training and order them into two distinct comma-separated sentences that start with “A video of”. We extract visual $A_{i}^{v}$ and textual $A_{i}^{t}$ coarse tag features by using the same text encoder used for the caption. Therefore, given a video/caption pair $(v_{i},t_{i})$ we get a quadruple $(V_{i},T_{i},A_{i}^{v},A_{i}^{t})$. Inspired by [3], we disentangle each element of the quadruple into $K$ independent, equal-sized latent concepts. For example, when disentangling $V_{i}$, we get $K$ independent latent concepts, i.e. $E_{i}^{v}=[e^{v}_{i,1},...,e^{v}_{i,K}]$. Each latent concept $e^{v}_{i,k}\in R^{d/K}$ represents a distinct concept and the independence of these factors ensures that each concept is uncorrelated to the other $K\text{-}1$ latent concepts, and is thus calculated by independently projecting the text representation:
$e^{v}_{i,k}=W^{v}_{k}V_{i}$ (2)
where $W^{v}_{k}$ is a trainable parameter. Similarly, $E_{i}^{t}$, $E_{i}^{a^{v}}$ and $E_{i}^{a^{t}}$ represent the latent concepts of the text $T_{i}$, the visual $A_{i}^{v}$ and the textual $A_{i}^{t}$ tag representations and are calculated in the same way. We name the $K$ latent concepts of the tags representation as the auxiliary visual concepts $E_{i}^{a^{v}}$ and auxiliary textual concepts $E_{i}^{a^{t}}$ respectively. We now have four disentangled representations for visual $E_{i}^{v}$, textual $E_{i}^{t}$, auxiliary visual $E_{i}^{a^{v}}$, and auxiliary textual $E_{i}^{a^{t}}$ concepts. Until now, these subspaces have been disentangled independently. We next describe how the alignment of these latent concepts can be used for enhancing cross-modal retrieval.
#Unique Visual Tags (VT) #Unique Textual Tags (TT)
Datasets #Videos # Captions Train Val Test Train Val Test
MSR-VTT-9k [5] 10,000 190,000 63,383 - 12,118 320,351 - 8,326
MSR-VTT-7k [70] 8,010 141,200 52,309 - 12,118 270,367 - 8,326
DiDeMo [71] 10,642 10,642 50,712 10,924 10,636 34,662 9,234 8,266
TGIF [72] 100,551 100,551 105,895 28,500 29,538 176,353 44,053 45,417
Charades [73] 9,848 9,848 41,538 - 15,360 25,406 - 9,874
YouCook2 [74] 13,829 13,829 33,786 - 16,457 26,035 - 12,227
TABLE I: Statistical analysis of tags after extraction.
Alignment of Disentangled Latent Concepts
@@
By default, approaches such as [3] directly align latent representations of videos and captions through a contrastive loss. Here, we extend this by applying the contrastive loss to align modality-specific concepts to the corresponding auxiliary concepts. In this way, modality-specific concepts assist MAC-VR in better aligning the video and text modalities. Specifically, we consider the visual concepts $E_{i}^{v}$ and the auxiliary visual concepts $E_{i}^{a^{v}}$. For each disentangled concept pair $(e_{i,k}^{v},e_{i,k}^{a^{v}})$ we minimise the distance between this pair then maximise the distance to other disentangled concepts, i.e. $(e_{i,k}^{v},e_{i,l}^{a^{v}});l\neq k$ in a contrastive fashion to align modality concepts. Recall that these latent concepts are learnt, and thus through this alignment, we aim to learn a representation of the video that matches the latent representations of the tags extracted from the VLM.
Similarly, we align the auxiliary textual concepts $E_{i}^{a^{t}}$ to the latent concepts $E_{i}^{t}$ extracted directly from the caption. We combine both modalities’ alignment of latent concepts to auxiliary latent concepts and refer to this as the Alignment Loss $\mathcal{L}_{A}$ which aligns $E_{i}^{v}$ with $E_{i}^{a^{v}}$ and $E_{i}^{t}$ with $E_{i}^{a^{t}}$.
Weighted Similarity and Training Loss
@@
Only a subset of visual concepts are usually described in the corresponding text. Therefore, we cannot directly leverage correlations between their latent concepts, so we use adaptive pooling to define weights for the visual and textual concepts and reduce their impact on the final similarity calculation. To do this, we design an adaptive module to estimate the confidence of each cross-modal concept matching. For each concept $k$, we concatenate the modality and auxiliary modality concepts $[e^{v}_{i,k},e^{t}_{i,k},e^{a^{v}}_{i,k},e^{a^{t}}_{i,k}]$ and use them to calculate the confidence of each cross-modal concept matching. We thus calculate:
$c_{i,k}=MLP([e^{v}_{i,k},e^{t}_{i,k},e^{a^{v}}_{i,k},e^{a^{t}}_{i,k}])$ (3)
If the confidence $c_{i,k}$ is low, the corresponding latent concept of the $k^{th}$ subspace is matched with a low score. Given this confidence, we aggregate all the visual and textual latent concept pairs to calculate the overall similarity of the video and text $S(v_{i},t_{i})$, through adaptive pooling as:
@@
Following common approaches, we use the InfoNCE loss [75, 76] as our Cross-Modal Loss ($\mathcal{L}_{S}$) to optimise the cross-modal similarity $S(v_{i},t_{i})$; the contrastive loss $\mathcal{L}_{C}$ to align the latent modality concepts (as introduced in [3]); and our proposed Alignment Loss $\mathcal{L}_{A}$ to align the modality with the auxiliary modality concepts:
$\begin{split}\mathcal{L}=\mathcal{L}_{S}(S(v_{i},t_{i}))+\alpha_{1}\mathcal{L}_{C}(E_{i}^{v},E_{i}^{t})+\\ \alpha_{2}\mathcal{L}_{A}(E_{i}^{v},E_{i}^{t},E^{a^{v}}_{i},E^{a^{t}}_{i}).\end{split}$ (5)
where $\alpha_{1}$ and $\alpha_{2}$ are weight parameters. During inference, the weighted similarity $S(v_{i},t_{i})$ as in Eq. 4 is used, given the query caption and every video in the gallery.
@@
Figure 4: Comparison of $RSum$ against baseline MAC-VR. $BS$: Batch Size. $N_{v}$: Number of Frames.

### v3

@@ 3.2 Tag Extraction
To extract visual and textual tags, we use a custom prompt (see Sec. C in supp.) to query the most general tags for the input video $v_{i}$ and caption $t_{i}$. We extract tags individually in an offline manner from both modalities so they can be used for training and inference. We call these tags as modality-specific tags because they are extracted from a single modality.
3.3 Architecture
@@
We start from a standard Text-Conditioned Video Encoder (T-CVE) before incorporating our proposed tags into each modalities’ latent concepts. These concepts are aligned and pooled to find the similarity between a video and a caption. Our architecture is shown in Fig. 3.
3.3.1 Text-Conditioned Video Encoder (T-CVE)
@@
Given a caption $t_{i}$, we extract its text representation $T_{i}\in\mathbb{R}^{d}$}. For the video representation, we first sample $N_{v}$ frames from a video $v_{i}$, then encode them and aggregate the embedding of all frames to obtain the frame representation $F_{j}$ with $j\in\{1,...,N_{v}\}$. Since captions often describe specific moments, as shown in previous works [Bain et al.(2022)Bain, Nagrani, Varol, and Zisserman, Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen, Gorti et al.(2022)Gorti, Vouitsis, Ma, Golestan, Volkovs, Garg, and Yu], matching only the relevant frames improves semantic precision and reduces noise. To achieve this, we aggregate the frame representations conditioned on the text. Firstly, we calculate the inner product between the text and the frame representation $F_{j}$ with $j\in\{1,...,N_{v}\}$:
@@ 3.3.2 Latent Concepts
To utilise the visual and textual tags extracted from foundational models in Sec 3.2, we first randomly pick $N$ visual and textual tags (per modality) during training and order them into two distinct comma-separated sentences that start with “A video of”. We extract visual $A_{i}^{v}$ and textual $A_{i}^{t}$ coarse tag features by using the same text encoder used for the caption. Therefore, given a video/caption pair $(v_{i},t_{i})$ we get a quadruple $(V_{i},T_{i},A_{i}^{v},A_{i}^{t})$. Inspired by [Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen], we disentangle each element of the quadruple into $K$ independent, equal-sized latent concepts. For example, when disentangling $V_{i}$, we get $K$ independent latent concepts, i.e. $E_{i}^{v}=[e^{v}_{i,1},...,e^{v}_{i,K}]$. Each latent concept $e^{v}_{i,k}\in R^{d/K}$ represents a distinct concept and the independence of these factors ensures that each concept is uncorrelated to the other $K\text{-}1$ latent concepts, and is thus defined by independently projecting the video representation as $e^{v}_{i,k}=W^{v}_{k}V_{i}$, where $W^{v}_{k}$ is a trainable parameter. Similarly, $E_{i}^{t}$, $E_{i}^{a^{v}}$ and $E_{i}^{a^{t}}$ represent the latent concepts of the text $T_{i}$, the visual $A_{i}^{v}$ and the textual $A_{i}^{t}$ tag representations, which are calculated in the same way. We name the $K$ latent concepts of the tags representation as the auxiliary visual concepts $E_{i}^{a^{v}}$ and auxiliary textual concepts $E_{i}^{a^{t}}$ respectively. We now have four disentangled representations for visual $E_{i}^{v}$, textual $E_{i}^{t}$, auxiliary visual $E_{i}^{a^{v}}$, and auxiliary textual $E_{i}^{a^{t}}$ concepts. Until now, these subspaces have been disentangled independently. We next describe how to align these latent concepts.
3.3.3 Alignment of Disentangled Latent Concepts
@@
By default, approaches such as [Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen] directly align latent representations of videos and captions through a contrastive loss. Here, we extend this by applying the contrastive loss to align modality-specific concepts to the corresponding auxiliary concepts. In this way, modality-specific concepts assist MAC-VR in better aligning the video and text modalities. Specifically, we consider the visual concepts $E_{i}^{v}$ and the auxiliary visual concepts $E_{i}^{a^{v}}$. For each disentangled concept pair $(e_{i,k}^{v},e_{i,k}^{a^{v}})$ we minimise the distance between this pair then maximise the distance to other disentangled concepts, i.e. $(e_{i,k}^{v},e_{i,l}^{a^{v}}):\,l\neq k$ in a contrastive fashion to align modality concepts. Recall that these latent concepts are learnt, and thus through this alignment, we aim to learn a representation of the video that matches the latent representations of the tags extracted from the VLM. Similarly, we align the auxiliary textual concepts $E_{i}^{a^{t}}$ to the latent concepts $E_{i}^{t}$ extracted directly from the caption. We combine both modalities’ alignment of latent concepts to auxiliary latent concepts and refer to this as the Alignment Loss $\mathcal{L}_{A}$ which aligns $E_{i}^{v}$ with $E_{i}^{a^{v}}$ and $E_{i}^{t}$ with $E_{i}^{a^{t}}$.
3.3.4 Weighted Similarity and Training Loss
@@
Only a subset of visual concepts are usually described in the corresponding text. Therefore, we cannot directly leverage correlations between their latent concepts, so we use adaptive pooling to define weights for the visual and textual concepts and reduce their impact on the final similarity calculation. To do this, we design an adaptive module to estimate the confidence of each cross-modal concept matching. For each concept $k$, we concatenate the modality and auxiliary modality concepts $e^{v}_{i,k},e^{t}_{i,k},e^{a^{v}}_{i,k},e^{a^{t}}_{i,k}]$ and use them to calculate the confidence of each cross-modal concept matching $c_{i,k}=MLP([e^{v}_{i,k},e^{t}_{i,k},e^{a^{v}}_{i,k},e^{a^{t}}_{i,k}])$. If the confidence $c_{i,k}$ is low, the corresponding latent concept of the $k^{th}$ subspace is matched with a low score. Given this confidence, we aggregate visual and textual latent concept pairs to calculate the overall similarity of the video and text $S(v_{i},t_{i})$, through adaptive pooling as:
@@
Following common approaches, we use the InfoNCE loss [Gutmann and Hyvärinen(2012), Józefowicz et al.(2016)Józefowicz, Vinyals, Schuster, Shazeer, and Wu] as our Cross-Modal Loss ($\mathcal{L}_{S}$) to optimise the cross-modal similarity $S(v_{i},t_{i})$; the contrastive loss $\mathcal{L}_{C}$ to align the latent modality concepts (as introduced in [Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen]); and our proposed Alignment Loss $\mathcal{L}_{A}$ to align the modality with the auxiliary modality concepts. We define our total loss as $\mathcal{L}=\mathcal{L}_{S}(S(v_{i},t_{i}))+\alpha_{C}\mathcal{L}_{C}(E_{i}^{v},E_{i}^{t})+\alpha_{A}\mathcal{L}_{A}(E_{i}^{v},E_{i}^{t},E^{a^{v}}_{i},E^{a^{t}}_{i})$, where $\alpha_{C}$ and $\alpha_{A}$ are hyperparameters to weigh the losses. During inference, the weighted similarity $S(v_{i},t_{i})$ as in Eq. 2 is used, given the query caption and every video in the gallery.
4 Experiments
4.1 Datasets, Implementation Details and Metrics
Datasets: We perform experiments on five video datasets: MSR-VTT [Xu et al.(2016)Xu, Mei, Yao, and Rui], DiDeMo [Hendricks et al.(2017)Hendricks, Wang, Shechtman, Sivic, Darrell, and Russell], TGIF [Li et al.(2016)Li, Song, Cao, Tetreault, Goldberg, Jaimes, and Luo], Charades [Sigurdsson et al.(2016)Sigurdsson, Varol, Wang, Farhadi, Laptev, and Gupta] and YouCook2 [Zhou et al.(2018)Zhou, Xu, and Corso], more details in Sec. D of the supp.
Architecture Details: We build our model using DiCoSA [Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen] as our baseline, employing CLIP’s VIT-B/32 [Radford et al.(2021b)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.] as the image encoder and CLIP’s transformer base as the text encoder to encode captions and visual/textual tags. Unless specified, we set $\tau_{a}=3$ in Eq. 1, $K=8$, and weight parameters $\alpha_{C}=\alpha_{A}=1$. Our MLP consists of two 256-length linear layers with a ReLU activation, and we employ a 4-layer transformer to aggregate frame embeddings.
Tag Extraction: We use the fine-tuned VideoLLaMA2 [Cheng et al.(2024b)Cheng, Leng, Zhang, Xin, Li, Chen, Zhu, Zhang, Luo, Zhao, et al.] for visual tags and Llama3.1 [Dubey et al.(2024)Dubey, Jauhri, Pandey, Kadian, Al-Dahle, Letman, Mathur, Schelten, Yang, Fan, Goyal, Hartshorn, Yang, Mitra, Sravankumar, Korenev, Hinsvark, Rao, Zhang, Rodriguez, Gregerson, Spataru, Rozière, Biron, Tang, Chern, Caucheteux, Nayak, Bi, Marra, McConnell, Keller, Touret, Wu, Wong, Ferrer, Nikolaidis, Allonsius, Song, Pintz, Livshits, Esiobu, Choudhary, Mahajan, Garcia-Olano, Perino, Hupkes, Lakomkin, AlBadawy, Lobanova, Dinan, Smith, Radenovic, Zhang, Synnaeve, Lee, Anderson, Nail, Mialon, Pang, Cucurell, Nguyen, Korevaar, Xu, Touvron, Zarov, Ibarra, Kloumann, Misra, Evtimov, Copet, Lee, Geffert, Vranes, Park, Mahadeokar, Shah, van der Linde, Billock, Hong, Lee, Fu, Chi, Huang, Liu, Wang, Yu, Bitton, Spisak, Park, Rocca, Johnstun, Saxe, Jia, Alwala, Upasani, Plawiak, Li, Heafield, Stone, and et al.] for textual tags. VideoLLaMA2 runs on $8$ sparsely sampled frames with temperature values $\tau\in{0.7,0.8,0.9,1.0}$ – the same values are used for Llama3.1 – to yield diverse tags. During training, we randomly select $N$ tags, while during inference we use fixed $M$ tags for deterministic results; we set $N=M=12$ for training and evaluating MAC-VR, except when using DSL, where $N=6$ and $M=8$. We extract modality-specific tags from all videos and captions in the datasets above. On average, we have 27272929 visual and textual tags per video/caption across all the datasets and splits. More tag extraction details in supp.
Train/Val Details: Following [Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen], we optimise the model using an Adam optimiser with linear warm-up, setting initial learning rates of 1e-7 for the text and video encoders and 1e-3 for other modules. We use $N_{v}=12$ frames and $BS=10$ for all datasets except DiDeMo, where $N_{v}=50$ for video-paragraph retrieval and $BS=64$ due to GPU memory.
Metrics: We present the retrieval performance for text-to-video retrieval task using standard metrics: Recall at $L=1,5$ ($R@L$) and mean rank ($MeanR$). We report the complete tables including $R@10$ and median rank $MR$ in Sec. E in the supp. We use the sum of cross-modal $R@L$ with $L=1,5,10$ as $RSum$ to rank methods.
4.2 Results
First, we present the comparison of MAC-VR against the baseline DiCoSA [Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen].
@@ 4.2 Results
Figure 4: Comparison of $RSum$ with baseline. BSBS: Batch Size. $N_{v}$: Number of Frames.
Dataset DiCoSA (s/sample) MAC-VR (s/sample) $\Delta$ Time (s/sample)
Charades 0.023 $\pm$ 0.001 0.057 $\pm$ 0.002 +0.034
DiDeMo 0.143 $\pm$ 0.003 0.179 $\pm$ 0.004 +0.036
MSRVTT 0.018 $\pm$ 0.001 0.065 $\pm$ 0.006 +0.047
TGIF 0.055 $\pm$ 0.001 0.077 $\pm$ 0.002 +0.022
YC2 0.020 $\pm$ 0.003 0.061 $\pm$ 0.001 +0.042
Table 1: Inference time comparison between DiCoSA and MAC-VR per sample.
Dataset VT (s/sample) TT (s/sample)
Charades 0.918 $\pm$ 0.103 0.964 $\pm$ 0.147
DiDeMo 0.927 $\pm$ 0.091 0.981 $\pm$ 0.021
MSRVTT 0.971 $\pm$ 0.319 0.941 $\pm$ 0.186
TGIF 0.555 $\pm$ 0.100 0.926 $\pm$ 0.116
YC2 1.043 $\pm$ 0.131 1.016 $\pm$ 0.141
Table 2: Visual and textual tag extraction times per sample for a single $\tau$.
Then, we compare MAC-VR with SOTA methods, particularly highlighting the impact of inference strategies on fairness of comparison.
Comparison with Baseline. In Fig. 4, we compare MAC-VR with our Baseline DiCoSA trained using the same training parameters, i.e. batch size, $BS$; same number of frames, $N_{v}$; and when using all inference strategies (IS). In all the datasets we improve $\Delta RSum$ compared to our baseline DiCoSA across all IS – +0.9+0.9 for MSR-VTT-9k, +3.4+3.4 for MSR-VTT-7k, +2.8+2.8 for DiDeMo, +9.5+9.5 for Charades, +3.0+3.0 for TGIF and +0.6+0.6 for YouCook2 when using DSL. Tab.2 presents the per-sample inference time comparison between DiCoSA[Jin et al.(2023b)Jin, Li, Cheng, Huang, Wang, Yuan, Liu, and Chen] and MAC-VR across all datasets, indicating that while MAC-VR increases its inference time slightly, it remains comparable to the original DiCoSA. Furthermore, Tab. 2 reports the extraction time of visual and textual tags for a single $\tau$ across datasets, which should be considered as an additional component of the total inference time.
Comparison with SOTA. In Tab. 3, we provide a comprehensive comparison of MAC-VR against all SOTA works. To ensure fairness, we divide methods by inference strategy (IS) and evaluate three settings: MAC-VR without IS, with QB, and with DSL. We exclude methods that use additional input modalities or stronger pre-training. For instance, in [Ibrahimi et al.(2023)Ibrahimi, Sun, Wang, Garg, Sanan, and Omar, Gabeur et al.(2022)Gabeur, Nagrani, Sun, Alahari, and Schmid, Chen et al.(2023a)Chen, Li, Wang, Zhao, Sun, Zhu, and Liu, Lin et al.(2022)Lin, Lei, Bansal, and Bertasius], models also train on audio, so they are not directly comparable to MAC-VR, which only uses video and text. Moreover, VAST [Chen et al.(2023a)Chen, Li, Wang, Zhao, Sun, Zhu, and Liu] includes subtitles (when available) in training as additional modality and is pre-trained on a multi-modal caption dataset. Similarly, CLIP-ViP [Xue et al.(2023)Xue, Sun, Liu, Fu, Song, Li, and Luo] uses extra pre-training datasets to improve retrieval performance further. We also omit T-MASS [Wang et al.(2024a)Wang, Sun, Wang, Liu, Dianat, Rabbani, Rao, and Tao] as the official code has been retracted due to data leakage. Overall, MAC-VR achieves best overall performance when combined with DSL across all the datasets. Results are consistent using QB for MSR-VTT (both splits) and without inference strategies on TGIF, Charades and YouCook2. On MSR-VTT (both splits), results are comparable to SOTA without IS. Across 5 of the 6 datasets, MAC-VR shows the best performance for any inference strategy. However, on DiDeMo, its performance is lower than SOTA methods for all inference strategies, even though the DSL approach still gives good results. This difference may be influenced by different training parameters (i.e. $BS$ and $N_{v}$) used in other SOTA [Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen, Guan et al.(2023)Guan, Pei, Shao, Liu, Li, Gu, Xu, Xu, Yan, and Lam, Jin et al.(2023a)Jin, Huang, Xiong, Tian, Liu, Ji, Yuan, and Chen, Jin et al.(2023c)Jin, Li, Cheng, Li, Ji, Liu, Yuan, and Chen, Wu et al.(2023a)Wu, Luo, Fang, Wang, and Ouyang, Wang et al.(2023b)Wang, Sung, Cheng, Bertasius, and Bansal, Li et al.(2023a)Li, Song, Gao, Zhu, and Shen, Chen et al.(2023b)Chen, Wang, Lin, Qi, Ma, and Shan]. To showcase this, we re-ran two SOTA methods [Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen, Wang et al.(2023b)Wang, Sung, Cheng, Bertasius, and Bansal] using our training settings, and both performed worse than in their original experiments, i.e. -$\text{-}6.4$ for [Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen] and -$\text{-}15.3$ for [Wang et al.(2023b)Wang, Sung, Cheng, Bertasius, and Bansal], and worse than our MAC-VR, i.e. -$\text{-}0.6$ for [Jin et al.(2022)Jin, Huang, Liu, Wu, Ge, Song, Clifton, and Chen] and -$\text{-}8.0$ for [Wang et al.(2023b)Wang, Sung, Cheng, Bertasius, and Bansal].